A stone is projected vertically upwards from a point on the ground with a velocity of 20 meter per second .tow seconds later another stone is projected from the same point with the same velocity .when and where do they meet given g=10 meter per second square

Grade:11

1 Answers

Shailendra Kumar Sharma
6 years ago
the first stone was sent at t=0 and second stone was thrown at t= 2 sec. not distance travelled by first one in 2 secs will be 
s=ut-1/2gt2
s=20*2-(.5)*10*4=20m  and velocity at this time will be v= u-at=20-2*10 =0 
Now after two seconds when second stone was thrown at 20m/s, the stones will have relative velocity of these particles will be 20m/s. a distace of 20 meter is to be covered with 20 m/s so the time taken will be 1 sec